OpenStudy (anonymous):

i need help solve each equation. check your solution y-6/y=5

OpenStudy (anonymous):

5/3p+2/3=5+p/2p

OpenStudy (anonymous):

\[\frac{y-6}{y}=5\] \[y-6=5y\] \[-6=4y\] \[y=-\frac{3}{2}\]